XxAnimeproxX:

Given f(x) = the quantity of 3x minus 1, divided by 2, solve for f−1(4). (5 points)

Cupcake123456:

x2 - 3x + 1 = fx x2 - 3x = -1 x2 - 3x + 9/4 = 5/4 (x-3/2)^2 - 5/4 = fx (answer)

Cupcake123456:

sooo.. work out the last step

chaanah:

shadow can u help me?

Shadow:

How did you get x^2 cupcake

Cupcake123456:

x^2 - 3*x + a million = 0 ,divide the two section with a million Which bring about x^2 - 3*x + a million = 0 , And the coefficient of x is -3 we could desire to continually use the undeniable fact that ( x + q )^2 = x^2 + 2*q*x + q^2 , and anticipate that q = -3/2 = -a million.5 meaning we are able to turn the equation into x^2 - 3*x + 2.25 - a million.25 = 0 And it particularly is the comparable with ( x - a million.5 )^2 - a million.25 = 0 the respond is C @shadow ?

Shadow:

\[f(x) = \frac{ 3x - 1 }{ 2 }\] Two simple steps 1. Inverse the function 2. Input 4 for x then solve for f(x) aka y

Shadow:

@XxAnimeproxX Do you know how to inverse functions?